OpenStudy (anonymous):

if f(x)=2x-4, find f(q +1) A) 2q + 4 B) 2q + 2 C) 2q - 6 D) 2q - 2

OpenStudy (anonymous):

Answer:B Just thing of g and f as numbers, g - f = then substitute in what the problem tells you g and f are, [(x^2) - 2x - 4] - [4x + 1] distribute the negative, (x^2) - 2x -4 - 4x -1 [then collect like terms] x^2 - 6x -5 so (g - f) (x) = x^2 - 6x -5

OpenStudy (anonymous):

oh ok